Question 4: (a) A robot moves on a NxN grid, NxN starting in the top-left corner S and map: moving either right or down at each step, until destination D is reached Robot Determine the big o complexity of a program that prints all possible paths from S to D. (2 marks) (b) Let A[1..nl be an array of bits (1 or 0 and F0 a function of time complexity 0(n). Give the time complexity of the following C++ code fragment. Justify your result by considering different cases for A. (2 marks) counter 0; for (i 1; i n i++) if (Ali] 1) counter++ else F(counter); counter 0;

data structure
Answer #1

Question 4 a)

A path from S to D will contain N-1 right moves and N-1 down moves. So, number of paths from S to D is equal to number of ways of arranging N-1 right moves and N-1 down moves

So, there are \frac{(N-1 + N-1)!}{(N-1)!(N-1)!} possible paths from S to D.

So time complexity to print all the possible paths = O(\frac{(N-1 + N-1)!}{(N-1)!(N-1)!})

Question 4 b)

Let m = number of zeros in array A.

So, Time complexity of given code = (m)*\theta(n) + (n-m)*(1)

= \theta(nm)